OpenStudy (anonymous):

can some check my work plz? What is the sum of a 6-term geometric sequence if the first term is 11, the last term is –11,264 and the common ratio is –4? The answer I got waz 9009? is this right?

OpenStudy (anonymous):

can u help me?

ganeshie8 (ganeshie8):

im getting -9009

OpenStudy (anonymous):

can you explain why plz? :)

ganeshie8 (ganeshie8):

ya sure.. .first can you show ur steps. which formula u used

OpenStudy (anonymous):

okay. I used sn=a1 1-rn/1-r is the formula I used , and then I replaced 11=a1 , -4=r the 6=n

OpenStudy (anonymous):

11*1-(-4)^6/1-(-4)

ganeshie8 (ganeshie8):

ok. so far it looks good

OpenStudy (anonymous):

then I subtracted the one from what I got (-4)^6 which is =4096/0ver 5

OpenStudy (anonymous):

the 5 is from 1-(-4)

ganeshie8 (ganeshie8):

yep wait a sec

OpenStudy (anonymous):

okay.

ganeshie8 (ganeshie8):

so you got this : 11 * (1-4096)/5

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yes.

ganeshie8 (ganeshie8):

continue... next step

OpenStudy (anonymous):

then I 11*4095/5 which = 45045/5 which =9009?

ganeshie8 (ganeshie8):

how is 1-4096 = 4095 ?

OpenStudy (anonymous):

huh?

ganeshie8 (ganeshie8):

it should be -4095 see ?

OpenStudy (anonymous):

and after I do that, I solve like I just did?

ganeshie8 (ganeshie8):

yes. thats the only sign part you did wrong.. rest everything is fine :)

OpenStudy (anonymous):

lol, I'm an idiot. Thanks :D

ganeshie8 (ganeshie8):

lol.. you did very good :DD

OpenStudy (anonymous):

thanks again you were a big help ;)

ganeshie8 (ganeshie8):

np... :)